When a stationary gas was compressed frictionless at constant pressure it's initial state of 12cm³ and 0.105mPa was found to change to final state of 120cm³ there was a transfer of 425kj of heat from the gas during the process.How much did the internal energy of the gas change


1
Expert's answer
2021-12-06T16:42:11-0500

"\Delta U=Q-A=Q-p\Delta V=414~kJ."